### v1

# Real-IAD Variety: Pushing Industrial Anomaly Detection Dataset to a Modern Era

Abstract
Industrial Anomaly Detection (IAD) is critical for enhancing operational safety, ensuring product quality, and optimizing manufacturing efficiency across global industries. However, the IAD algorithms are severely constrained by the limitations of existing public benchmarks. Current datasets exhibit restricted category diversity and insufficient scale, frequently resulting in metric saturation and limited model transferability to real-world scenarios. To address this gap, we introduce Real-IAD Variety, the largest and most diverse IAD benchmark, comprising 198,960 high-resolution images across 160 distinct object categories. Its diversity is ensured through comprehensive coverage of 28 industries, 24 material types, and 22 color variations Our comprehensive experimental analysis validates the benchmark’s substantial challenge: state-of-the-art multi-class unsupervised anomaly detection methods experience significant performance degradation when scaled from 30 to 160 categories. Crucially, we demonstrate that vision-language models exhibit remarkable robustness to category scale-up, with minimal performance variation across different category counts, significantly enhancing generalization capabilities in diverse industrial contexts. The unprecedented scale and complexity of Real-IAD Variety position it as an essential resource for training and evaluating next-generation foundation models for anomaly detection. By providing this comprehensive benchmark with rigorous evaluation protocols across multi-class unsupervised, multi-view, and zero-/few-shot settings, we aim to accelerate research beyond domain-specific constraints, enabling the development of scalable, general-purpose anomaly detection systems. Real-IAD Variety will be made publicly available to facilitate innovation in this critical field.
1 Introduction
The field of anomaly detection has witnessed a paradigm shift towards fully unsupervised [1, 2, 3], multi-class unsupervised (unified) [4, 5, 6], and zero-/few-shot [7, 8, 9] learning frameworks, marking a transformative era in model design. Within the specialized domain of Industrial Anomaly Detection (IAD), substantial progress has been achieved in addressing the inherent complexity and variability of industrial processes.
Figure 1: Industry distribution of the proposed Real-IAD Variety dataset. The dataset encompasses 8 major industrial categories (denoted as c): electrical, transport, cultural products, metal, general, electronics, rubber plastic, and other manufacturing sectors. These major categories are further subdivided into 28 industrial subcategories (denoted as s). Complete nomenclature details are provided in Appendix.
Recent IAD research has predominantly concentrated on developing unified models capable of detecting anomalies across multiple heterogeneous industrial systems or processes [4, 10, 11, 12, 5]. In parallel, zero-shot and few-shot anomaly detection approaches [13, 7, 8, 9] have emerged to identify anomalies in previously unseen domains without requiring domain-specific prior knowledge. These advances are fundamentally transforming IAD, enabling the development of more robust, efficient, and adaptable detection frameworks.
Despite the notable contributions of existing benchmarks such as MVTec AD [14], VisA [15], PAD [16], and Real-IAD [17], these datasets exhibit critical limitations in category diversity and scenario coverage, thereby constraining the comprehensive evaluation of contemporary unified and zero-/few-shot IAD models. Specifically, Real-IAD [17], which currently offers the largest category count, contains only 30 categories, while the widely adopted MVTec AD and VisA datasets comprise merely 15 and 12 categories, respectively. Although recent efforts [18, 19, 20] have attempted to synthesize large-scale anomaly images using pre-trained diffusion models, these approaches continue to face challenges in generating authentic and diverse anomalies.
To advance IAD research toward unified and zero-/few-shot paradigms, there exists an urgent need for datasets that encompass substantially broader industrial categories and defect typologies. To address these limitations and propel IAD datasets into a new era, we introduce Real-IAD Variety, a large-scale benchmark that represents a significant advancement in the field. As illustrated in Figure 1 and Figure 3, this dataset substantially expands the scope of IAD data collection, featuring 160 categories spanning 28 industries, 24 material types, and 22 color variations, with a total of 198,960 images accompanied by pixel-level annotations. The category count of Real-IAD Variety is approximately 5.3 times that of Real-IAD [17], the previously largest IAD dataset. By dramatically increasing defect type diversity, Real-IAD Variety becomes the first multi-industry IAD dataset, thereby overcoming the inherent constraints of prior benchmarks for unified and zero-/few-shot evaluation.
Real-IAD Variety provides comprehensive coverage across 28 industries, including electrical manufacturing, transportation equipment, and cultural products, addressing the industry diversity gap present in Real-IAD, which primarily focuses on multi-view detection challenges. In contrast, existing datasets such as VisA [15] and MVTec AD [14] predominantly target narrower industrial sectors. Furthermore, Real-IAD Variety is designed to reflect real-world industrial requirements by incorporating 24 material types—including composite materials, metals, plastics, wood, and ceramics—offering substantially broader and more practical material representation compared to the limited ranges in Real-IAD [17], VisA [15], and MVTec AD [14].
The introduction of Real-IAD Variety not only augments training data diversity but also enables systematic investigation of the generalization capabilities of state-of-the-art (SOTA) IAD methods. We establish comprehensive benchmarks across three critical IAD settings: (1) multi-class unsupervised anomaly detection, (2) multi-view anomaly detection, and (3) zero-/few-shot anomaly detection. These settings are designed to reflect the diverse challenges encountered in real-world IAD applications, providing a rigorous framework for performance evaluation. Through extensive experimentation, we identify two key findings:
First, a substantial increase in category count leads to significant performance degradation in multi-class unsupervised anomaly detection (MUAD) methods, revealing that current MUAD algorithms exhibit limited scalability with respect to category diversity. For instance, as demonstrated in Table 3, when the category count increases from 30 to 60, 100, and finally 160, existing algorithms experience a performance decline of 1030%.
Second, zero-/few-shot learning approaches demonstrate greater robustness to category count variations, and when the number of categories exceeds a certain threshold, their performance surpasses that of MUAD methods. As shown in Table 4, pre-trained vision-language zero-/few-shot anomaly detection models maintain nearly consistent performance regardless of category count.
In summary, our contributions are threefold:
  • We present Real-IAD Variety, a large-scale IAD dataset featuring unprecedented category diversity, extensive industry coverage, and high-quality pixel-level annotations. It expands the category count to 160 across 28 industries, 24 material types, and 22 color variations with 198,960 images, representing a substantial advancement over existing IAD benchmarks.
  • We establish comprehensive benchmarks on Real-IAD Variety across three critical IAD settings: multi-class unsupervised anomaly detection, multi-view anomaly detection, and zero-/few-shot anomaly detection, facilitating systematic evaluation of current methodologies.
  • We reveal that increasing category count significantly impairs the performance of multi-class and multi-view unsupervised anomaly detection methods, while zero-/few-shot approaches exhibit minimal sensitivity to category variations, suggesting a fundamental shift in scalability characteristics across different learning paradigms.
2 Related Work
2.1 IAD Datasets
In Industrial Anomaly Detection (IAD), dataset diversity and scale are critical determinants of model effectiveness. Current 2D IAD benchmarks, including MVTec AD [14], VisA [15], MVTec Caption [21], and Real-IAD [17], have demonstrated considerable success in establishing foundational evaluation frameworks. Real-IAD [17], the largest existing dataset, encompasses 30 categories, while MVTec AD [14] and VisA [15] contain 15 and 12 categories, respectively. The MVTec Caption dataset [21] augments MVTec AD with descriptive textual annotations; however, it remains confined to a limited set of object types. Additional datasets such as MTD [22], MPDD [23], BTAD [24], and KolektorSDD [25] contribute valuable resources but are constrained by their limited scale and category diversity. Despite these contributions, existing datasets exhibit insufficient defect category variability and industrial domain coverage, thereby limiting their utility for evaluating unified and zero-/few-shot IAD paradigms. In contrast, our proposed Real-IAD Variety addresses these limitations by substantially enhancing category diversity, industrial scope, and defect typology, thereby advancing the field toward more comprehensive and realistic evaluation scenarios.
2.2 Unsupervised Anomaly Detection
Traditional unsupervised anomaly detection (UAD) algorithms have predominantly focused on single-class, single-view 2D images, training separate models for each object category using exclusively normal samples. These methods are typically categorized into three paradigms: (1) Discriminative methods (e.g., CutPaste [26], DRAEM [27], and SimpleNet [11]), which learn decision boundaries by distinguishing normal data from synthetically generated anomalies; (2) Reconstruction-based methods (e.g., DAE [28], O-CGAN [29], and RD [30]), which operate under the assumption that anomalous regions exhibit higher reconstruction errors compared to normal patterns; (3) Embedding-based methods (e.g., PatchCore [31], CFA [10], and CSFlow [32]), which model the distribution of normal feature representations in a frozen or learned embedding space. While these approaches have achieved notable success in single-class scenarios, their scalability to multi-class and cross-domain settings remains limited, motivating the development of unified detection frameworks.
2.3 Multi-class Unsupervised AD
To systematically evaluate the generalization capabilities of single-modality IAD models, the multi-class unsupervised anomaly detection (MUAD) setting was first introduced in UniAD [4]. This paradigm challenges models to train on all object categories within a dataset simultaneously, employing a single unified model for both training and inference. MUAD benchmarks encompass both traditional UAD methods, such as DRAEM [27], SimpleNet [11], CFA [10], CFLOW-AD [33], and RD [30]—and methods explicitly designed for unified training, including UniAD [4], OneNIP [5], MambaAD [12], DesTSeg [34], and Dinomaly [6]. However, existing MUAD evaluations have been conducted on datasets with limited category counts (typically $\leq$ 30 classes), leaving the scalability of these methods to large-scale, highly diverse industrial scenarios largely unexplored. Our Real-IAD Variety dataset, with its 160 categories, enables the first comprehensive assessment of MUAD performance under realistic industrial diversity.
2.4 Multi-view Anomaly Detection
Beyond single-image analysis, Multi-View Anomaly Detection (MVAD) leverages complementary information from multiple viewpoints to enhance inspection accuracy and robustness. This approach originates from the stringent precision requirements of industrial quality control, where anomalies may be occluded or imperceptible from a single perspective. Representative methods such as MVAD [35] exploit multi-view consistency constraints and cross-view feature fusion to improve detection performance, which is critical for minimizing false negatives in practical deployment scenarios. Nevertheless, existing MVAD datasets predominantly focus on limited object types and controlled imaging conditions, constraining the evaluation of cross-view generalization in diverse industrial contexts. Real-IAD Variety addresses this gap by providing multi-view annotations across 160 categories, facilitating rigorous MVAD benchmarking.
2.5 Zero-/Few-Shot AD with VLMs
Figure 2: Data collection and annotation pipeline for the proposed Real-IAD Variety. The pipeline comprises a four-stage sequential process: (a) Material Preparation: This initial phase encompasses the assembly of a diverse array of materials, spanning 160 distinct categories sourced from 28 industrial domains and encompassing 24 material compositions. (b) Acquisition Equipment Design: The second phase involves the design of data capture apparatus, comprising one $\text{top}-\text{down}$ camera for overhead views and four peripheral cameras to capture lateral perspectives. (c) Data Collection and Annotation: The third phase pertains to the data collection process, which includes meticulous pixel-level manual annotation, rigorous algorithmic cross-validation, and iterative refinement. This process iterates until the model’s predicted Average Precision (AP) scores exhibit negligible variation below a predetermined threshold, following the methodology established in Real-IAD [17]. (d) Defect Taxonomy: The lower section illustrates 23 distinct defect types alongside their characteristic visual representations. Zoom in for enhanced visibility of defect regions delineated in red.
The vision-language models (VLMs) have emerged as a promising direction for anomaly detection, enabling semantic reasoning and cross-domain generalization through natural language supervision. Recent advances have explored zero-shot (ZSAD) and few-shot (FSAD) anomaly detection using large-scale pre-trained models such as CLIP [36]. WinCLIP [37] pioneered the application of vision-language models for zero-shot anomaly classification and segmentation by introducing dual-class textual prompts and patch-wise window operations. AnomalyCLIP [7] employs object-agnostic textual prompts to capture generalizable normal and abnormal representations, further enhanced through auxiliary anomaly detection datasets. AdaCLIP [38] introduces learnable hybrid prompts and regional anomaly feature extraction to improve detection precision, while VCP-CLIP [39] designs Pre-VCP and Post-VCP modules to optimize textual embeddings with visual contextual prompts, promoting cross-modal information interaction.
Recently, AdaptCLIP [8] treats CLIP as a foundational service, proposing alternating and comparative learning strategies based on three lightweight adapters to support zero- and few-shot generalization across domains. Despite these methodological advances, existing zero-/few-shot approaches have been validated primarily on small-scale datasets with limited category diversity (typically $\leq$ 30 classes), leaving their scalability and robustness under large-scale, multi-industry conditions unverified. Our Real-IAD Variety enables the first large-scale evaluation of vision-language anomaly detection methods across 160 categories, thereby facilitating systematic investigation of their generalization capabilities in realistic industrial settings.
Figure 3: Statistical characteristics of Real-IAD Variety across multiple dimensions. (a) Anomalous region proportion: Real-IAD Variety exhibits a broader and more balanced distribution of anomalous region proportions relative to total image area compared to Real-IAD [17], substantially increasing dataset complexity. (b) Defect aspect ratio: Real-IAD Variety provides diverse aspect ratios for minimum bounding rectangles of defects, comparable to Real-IAD, introducing additional diversity and detection challenges. Representative samples are shown for intuitive visualization. (c) Material distribution: Real-IAD Variety encompasses 24 material types for practical applications, imposing higher requirements on method robustness. (d) Color distribution: Real-IAD Variety captures a wide color spectrum, which is essential for color-based anomaly detection research.
3 Methodology: Real-IAD Variety
3.1 Data Collection Pipeline
Inspired by the methodology established in Real-IAD [17], we construct a rigorous three-stage data collection pipeline to ensure dataset comprehensiveness and annotation quality, as illustrated in Figure 2. The pipeline encompasses material preparation, acquisition equipment design, and data collection with iterative annotation refinement.
Stage 1: Diverse Material Preparation. To construct the Real-IAD Variety dataset, a team of 12 members dedicated 11,000 working hours to material selection and procurement, simulating a wide spectrum of real-world defects based on material characteristics. To enhance dataset representativeness, we assembled 160 object categories spanning 24 material types across 28 industrial domains. Representative samples are depicted in Figure 2a. Given that industrial production typically achieves yield rates exceeding 99%, naturally occurring defective parts are scarce. Therefore, leveraging extensive production line experience, we artificially introduced four defect variations for each material category, aggregating to 23 major defect types across the entire dataset.
[…]
Stage 3: Data Collection and Annotation Refinement.
The data collection and annotation process for Real-IAD Variety adheres to the rigorous standards established by Real-IAD [17]. This stage involves meticulous pixel-level manual annotation, algorithmic cross-validation, and iterative refinement. The annotation process iterates until the model’s predicted Average Precision (AP) scores exhibit negligible variation below a predetermined threshold, ensuring annotation consistency and quality.
3.2 Comparison with Existing IAD Datasets
[…]
3.3 Dataset Characteristics and Statistical Analysis
Table 1: Comparison with popular 2D IAD datasets on different attributes. ✔: Satisfied. ✘: Unsatisfied.
Datasets Classes The Number of Images
Image
Resolution
Anomaly
Masks
Multiple
Views
Defect Types
Normal Anomaly All
[…]
MPDD [23] 6 1,064 282 1,346 $\sim1,024$ 8
MAD-Real [16] 10 540 221 761 $\sim3,472$ 2
MAD-Sim [16] 20 4,838 4,951 9,789 $\sim800$ 3
[…]
Statistical Distribution Analysis. Figure 3 illustrates the advantages and challenges of Real-IAD Variety compared to existing IAD datasets across multiple dimensions: (a) Anomalous region proportion: Compared to Real-IAD [17], Real-IAD Variety exhibits a broader and more uniform distribution of anomaly region proportions (Figure 3a), encompassing defects of varying sizes and thereby increasing detection difficulty. (b) Defect aspect ratio: The defect aspect ratio distribution in Real-IAD Variety is comparably diverse to Real-IAD (Figure 3b), presenting substantial challenges for models to discriminate defects of various geometric configurations. (c) Material diversity: Real-IAD Variety encompasses 24 material types (Figure 3c) across 28 industrial domains (Figure 1), providing comprehensive coverage from a practical application perspective. (d) Color spectrum: The dataset captures a wide color distribution (Figure 3d), which is critical for evaluating color-based anomaly detection approaches. These multifaceted characteristics substantially enhance dataset diversity and provide robust support for large-scale anomaly detection algorithm research.
Representative Defect Visualization. Figure 2d presents qualitative visualizations of 23 distinct defect types across various object categories. Notably, identical defect types manifest in diverse visual forms across different materials, reflecting the complexity of real-world industrial scenarios. These defect types, derived from authentic production line observations, enhance the practical applicability of Real-IAD Variety.
[…]
4 Experiments
4.1 Dataset Protocol and Evaluation Metrics
Training and Testing Protocol. The Real-IAD Variety dataset is partitioned into training and testing subsets, comprising 3,991 normal samples (19,955 images) for training and 35,799 samples (3,990 normal and 31,809 anomalous, totaling 178,995 images) for testing. Notably, the testing subset exhibits a nearly balanced distribution between normal and anomalous instances. To systematically evaluate the impact of category count on model performance, we partition Real-IAD Variety into three subsets: S1, S2, and S3, containing 30, 60, and 100 categories, respectively. The category selection employs randomization to mitigate potential biases in color and material representation.
Evaluation Metrics. Following established anomaly detection protocols, we employ three standard metrics: Image-level Area Under the Receiver Operating Characteristic Curve (I-AUROC) [27], Pixel-level AUROC (P-AUROC) [27], and Pixel-level Area Under the Precision-Recall Curve (P-AUPR) [15] to assess method effectiveness across classification and localization tasks.
4.2 Benchmark on Multi-Class Unsupervised AD
Experimental Setting. The Multi-Class Unsupervised Anomaly Detection (MUAD) paradigm, first introduced in UniAD [4], trains a unified model on all object categories within a dataset simultaneously, employing the same model for both training and inference. This approach eliminates the need for class-specific models, thereby reducing storage costs and potentially enabling the learning of generalizable features across multiple categories. This setting holds significant practical value for real-world industrial deployment.
Table 2: MUAD performance comparisons on Real-IAD Variety dataset. Dis., Emb., and Rec., respectively, represent discrimination-based, embedding-based, and reconstruction-based methods. Bold indicates the best results.
Methods I-AUROC P-AUROC P-AUPR
Dis.
DREAM [27] 49.6 52.1 1.3
SimpleNet [11] 54.5 70.6 3.6
Emb.
CFA [10] 52.3 56.8 2.2
CFLOW-AD [33] 69.8 86.5 18.0
Rec.
RD [30] 73.0 89.5 18.9
MabaAD [12] 79.5 90.8 28.2
UniAD [4] 64.2 86.4 11.3
DesTSeg [34] 76.2 67.4 35.2
Dinomaly [6] 81.4 91.5 37.6
Dinomaly+ 81.6 91.9 40.7
Results and Analysis under MUAD.
Table 3: MUAD performance comparisons across different anomaly detection datasets. Performance is measured using I-AUROC, P-AUROC, and P-AUPR metrics.
Datasets # Classes CFLOW-AD [33] SimpleNet [11] UniAD [4] MambaAD [12] Dinomaly [6] Dinomaly+
MVTec 15 80.4 90.7 37.1 78.2 81.0 24.8 96.5 96.8 44.7 98.6 97.7 56.3 99.6 98.4 69.3 99.7 98.7 76.9
ViSA 12 69.0 91.4 16.8 89.2 95.3 33.1 90.8 98.4 33.6 94.3 98.5 39.4 98.7 98.7 53.2 98.8 98.7 54.6
Real-IAD 30 55.7 81.3 1.6 57.2 75.7 2.8 83.0 97.3 21.1 86.3 98.5 33.0 89.3 98.8 42.8 90.0 98.9 47.6
Real-IAD Variety S1 30 55.0 63.8 2.9 75.0 84.8 21.7 79.9 89.7 24.4 87.0 92.3 39.4 90.1 92.4 54.1 90.5 92.5 57.0
Real-IAD Variety S2 60 54.2 60.0 2.6 67.1 82.0 13.3 66.9 86.7 14.1 85.8 92.3 37.8 88.1 92.4 49.0 88.7 92.7 53.1
Real-IAD Variety S3 100 52.9 57.6 2.4 59.3 74.7 5.7 66.7 86.9 12.9 85.0 92.3 35.7 85.4 92.2 43.8 85.7 92.3 47.2
Real-IAD Variety 160 52.3 56.8 2.2 54.5 70.6 3.6 64.2 86.4 11.3 79.5 90.8 28.2 81.4 91.5 37.6 81.6 91.7 42.4
Unless otherwise specified, MUAD models are trained at $\times256$ and a total of 100 epochs. For Dinomaly [6], we follow the original setting with $\times448$ resolution for 50,000 steps (e.g., approximately 160 epochs for Real-IAD Variety). Dinomaly+ extends Dinomaly [6] by incorporating an additional refined segmentation head [5], trained for 4 epochs on pseudo-anomalous data to enable coarse-to-fine anomaly localization.
Competitive Methods. We evaluate state-of-the-art unsupervised anomaly detection methods to comprehensively assess their scalability to large-scale datasets. Specifically, we benchmark discrimination-based methods (DRAEM [27], SimpleNet [11]), embedding-based methods (CFA [10], CFLOW-AD [33]), and reconstruction-based methods (RD [30], MambaAD [12], UniAD [4], DesTSeg [34], Dinomaly [6], and Dinomaly+) on Real-IAD Variety. Considering computational resource constraints, we focus on representative and strong baselines—CFLOW-AD [33], SimpleNet [11], UniAD [4], MambaAD [12], Dinomaly [6], and Dinomaly+ for comprehensive MUAD evaluation. Notably, PatchCore [31] is excluded from this comparison due to its memory bank mechanism, which incurs prohibitive memory consumption as category count increases, leading to potential out-of-memory issues on large-scale datasets like Real-IAD Variety.
Table 2 presents quantitative results of different methods on Real-IAD Variety. Several critical observations emerge that differ from findings on small-scale datasets:
All evaluated methods experience substantial and non-saturating performance degradation on Real-IAD Variety, indicating the dataset’s significant challenge for future research and confirming successful avoidance of the metric saturation phenomenon frequently observed in smaller-scale IAD benchmarks.
Discrimination-based approaches (DRAEM [27], SimpleNet [11]) and the embedding-based method CFA [10] exhibit near-failure performance, with I-AUROC scores approaching the 50% random guessing threshold. This demonstrates their severe lack of generalization capability when confronted with large-scale, high-variance normal data distributions.
Reconstruction-based methods maintain relatively satisfactory performance despite the challenging conditions. Notably, Dinomaly+ (building upon Dinomaly [6] and OneNIP [5]) achieves the best overall results across all metrics (I-AUROC: 81.6%, P-AUROC: 91.9%, P-AUPR: 40.7%), underscoring its superior generalization and robustness across large-scale multi-class scenarios. In contrast, UniAD [4], an early MUAD method, exhibits one of the most severe performance drops among reconstruction-based approaches.
Performance Trends with Increasing Categories. To evaluate performance trends as category count increases, we selected five representative methods across different paradigms. Table 3 presents results on several datasets and Real-IAD Variety with incrementally increasing categories: S1 (30), S2 (60), S3 (100), and the full 160 categories (visualized in Figure 4). Several consistent conclusions emerge under the MUAD setting:
Real-IAD Variety presents greater challenges compared to the original Real-IAD dataset. Despite maintaining equivalent category counts (30 classes), performance metrics on Real-IAD Variety are consistently lower, suggesting higher inherent complexity or greater intra-class variance.
A fundamental scalability challenge of MUAD is revealed: consistent performance degradation occurs across all evaluated methods as category count increases. Crucially, the magnitude of degradation is method-dependent, highlighting that different architectures possess varying levels of category generalization capability and robustness to scale-up.
Methods leveraging powerful backbones (e.g., DINOv2-R [41]) and advanced architectures (e.g., Transformer decoders), such as Dinomaly and Dinomaly+, consistently achieve substantial performance superiority across all datasets. This advantage is especially prominent on high-scale benchmarks, underscoring their effectiveness over normalizing flow-based (e.g., CFLOW-AD) or simpler backbone-based models (e.g., SimpleNet, UniAD, MambaAD).
Table 4: Comparisons of ZSAD and FSAD methods across different benchmarks. Performance is measured using I-AUROC, P-AUROC, and P-AUPR metrics. Delta indicates the performance gap between maximum and minimum values.
Datasets # Classes Zero-Shot 1-Shot 2-Shot
AnomalyCLIP [7] AdaCLIP [38] VCPCLIP [39] AdaptCLIP [8] AdaptCLIP [8] AdaptCLIP [8]
Real-IAD 30 69.5 94.8 26.6 71.8 80.3 34.6 71.7 95.9 29.2 74.2 94.9 28.2 81.8 97.1 36.6 82.9 97.3 37.8
Real-IAD Variety S1 30 67.7 87.9 33.7 73.5 73.4 34.8 74.0 88.6 33.6 73.2 88.1 36.7 79.4 90.7 49.0 81.2 91.0 50.4
Real-IAD Variety S2 60 69.0 88.2 33.5 74.0 74.3 37.7 76.1 89.1 36.0 72.6 88.1 36.5 78.7 91.0 48.0 80.4 91.3 49.8
Real-IAD Variety S3 100 70.0 89.1 33.8 75.2 75.1 38.5 76.1 90.2 36.5 72.8 89.1 36.8 79.2 91.8 48.3 81.0 92.0 50.2
Real-IAD Variety 160 69.8 89.3 32.9 74.4 74.5 38.9 75.9 90.6 36.2 73.0 89.2 36.2 79.7 92.0 48.0 81.4 92.2 49.7
Delta   $\downarrow$ 2.3 6.9 7.2 3.4 7.0 4.3 4.4 7.3 7.3 1.6 6.8 8.6 3.1 6.4 12.4 2.5 6.3 12.6
Figure 4: Performance trends of I-AUROC and P-AUPR metrics with increasing categories for MUAD, ZSAD and FSAD methods.
4.3 Benchmark on Zero-Shot and Few-Shot AD
Experimental Setting. Zero-Shot Anomaly Detection (ZSAD) requires models trained on auxiliary datasets to identify anomalies in novel object categories without class-specific training data. This setting is particularly valuable because obtaining comprehensive coverage of all object classes and anomaly types is impractical in real-world scenarios. ZSAD aims to enhance model robustness and adaptability by facilitating knowledge transfer from known to unknown categories, thereby enabling effective detection of novel anomalies across diverse and dynamic environments. To further evaluate model adaptability in open-world scenarios, we introduce a Few-Shot Anomaly Detection (FSAD) setting alongside ZSAD evaluation. In FSAD, a limited number of normal images ($k=\{1,2\}$) from the target class serve as visual prompts during inference, with their detection results integrated with ZSAD outputs to enhance final anomaly detection performance.
Competitive Methods. We evaluate AnomalyCLIP [7], AdaCLIP [38], VCP-CLIP [39], and AdaptCLIP [8] on Real-IAD Variety in zero-shot manner, conducting comprehensive comparisons with Real-IAD [17]. Given that Real-IAD Variety comprises 160 novel categories, it serves as an exceptionally challenging benchmark for assessing ZSAD model robustness and cross-category generalization in complex real-world scenarios. Notably, AdaptCLIP [8] supports both zero- and few-shot generalization across domains; therefore, we employ it to evaluate FSAD performance on Real-IAD Variety. Results and Analysis under ZSAD and FSAD. Table 4 presents ZSAD and FSAD results on Real-IAD and Real-IAD Variety, with performance trends visualized in Figure 4. Several key observations emerge:
ZSAD models demonstrate remarkable stability and robustness, while traditional MUAD methods exhibit significant performance degradation as category complexity increases. This fundamental difference is quantitatively highlighted by low I-AUROC delta values (e.g., 2.3 for AnomalyCLIP, 1.6 for AdaptCLIP) across increasing category counts (30 to 160). This suggests that VLM-based ZSAD approaches, which leverage external knowledge from vast pre-training data rather than modeling intrinsic training set distributions, are inherently less sensitive to data distribution shifts induced by category scale-up.
FSAD yields substantial performance gains using minimal class-specific prompt information. Specifically, using only two normal images (2-Shot), AdaptCLIP’s performance on Real-IAD Variety (160 classes) improves significantly, with I-AUROC rising from 73.0% (ZSAD) to 81.4% (2-Shot). This improvement confirms the high value of incorporating even minimal normal visual prompts to rapidly refine generic VLM knowledge for precise, target-specific anomaly discrimination. The consistent improvement from 1-Shot to 2-Shot further validates the effectiveness of this adaptation mechanism.
Few-shot AdaptCLIP (2-Shot) achieves the best overall performance across all benchmarks. On the most challenging 160-class Real-IAD Variety, 2-Shot AdaptCLIP reaches I-AUROC of 81.4%, which is highly competitive with the best MUAD model (Dinomaly+: 81.6%) that utilizes full normal training data. Crucially, delta analysis reveals that while few-shot adaptation boosts absolute performance, it incurs slightly higher delta for pixel-level metrics, suggesting that local adaptation mechanisms, while powerful, exhibit greater sensitivity to category complexity than global ZSAD paradigms.
The superior performance and inherent stability of VLM-based ZSAD/FSAD methods strongly advocate for multimodal knowledge integration. To further unlock the potential of these models, particularly in the complex cross-category setting of Real-IAD Variety, future research should focus on developing enhanced methods that effectively utilize detailed textual and semantic information, bridging the gap between high-level semantic understanding and fine-grained localized visual anomaly detection.
4.4 Benchmark on Multi-View AD
Experimental Setting. Multi-View Anomaly Detection (MVAD) originates from high-precision demands in industrial quality inspection. It aims to achieve detection capabilities unattainable from single-view images by integrating multiple perspective images of the same sample, thereby identifying defects that are invisible from certain viewpoints but visible from others. Additionally, this setting leverages multi-view information and consistency constraints to further enhance model performance.
Table 5: MVAD [35] performance on Real-IAD Variety with increasing categories.
Datasets # Classes I-AUROC P-AUROC P-AUPR
Real-IAD 30 86.6 97.9 30.3
Real-IAD Variety S1 30 85.5 91.9 34.7
Real-IAD Variety S2 60 82.9 91.6 29.9
Real-IAD Variety S3 100 81.8 91.5 26.8
Real-IAD Variety 160 78.1 90.4 22.3
Results and Analysis under MVAD. Table 5 presents quantitative results of MVAD [35] on Real-IAD Variety. As category count increases, model performance gradually decreases, consistent with trends observed in MUAD settings. In practical production line applications, multi-view imaging is essential to reduce false negative rates. Nevertheless, current metric results remain relatively modest, presenting a significant challenge for future research to address scalability and robustness in large-scale multi-view anomaly detection scenarios.
5 Future Work
Foundation Models for Anomaly Detection: Scaling Toward a New Paradigm. Foundation models have fundamentally transformed how diverse tasks are addressed through unified architectures, ushering in a new era of model design [42, 43, 44]. A critical enabler of this transformation is the availability of large-scale, high-quality datasets, which are essential for training versatile and generalizable models across various domains. In the context of Industrial Anomaly Detection (IAD), the development of foundation models is heavily dependent on comprehensive vision-language annotations. These annotations are indispensable not only for accurate anomaly detection but also for providing contextual understanding that enables models to reason about anomalies, thereby enhancing their generalization capabilities across diverse industrial scenarios.
Multimodal Anomaly Detection with Large-Scale Diverse Data. The introduction of Real-IAD Variety opens several promising research directions for advancing anomaly detection. First, the development of robust generalized anomaly detection models represents a critical avenue. Future research should explore training strategies that effectively integrate these textual insights with visual features, e.g., MMAD [45], thereby enhancing the robustness and adaptability of anomaly detection systems. Second, addressing logical anomalies in image-based anomaly detection remains an important challenge. Current methods often struggle with logical anomalies, where visual representations alone fail to capture complete contextual relationships among objects or components. Our preliminary ZSAD and FSAD experiments have revealed performance limitations in this regard. To address these constraints, we propose enabling text encoder fine-tuning to enhance model adaptability and contextual reasoning capabilities. Future research should focus on integrating textual descriptions into existing anomaly detection frameworks, enabling models to reason about logical relationships and contextual cues beyond purely visual patterns.
6 Conclusion
Industrial Anomaly Detection (IAD) is undergoing a fundamental transformation with the emergence of large-scale models capable of operating in unified and zero-/few-shot settings. This work makes a substantial contribution to the IAD dataset landscape by introducing Real-IAD Variety, a large-scale benchmark that not only expands the scope of available data but also establishes rigorous evaluation protocols for assessing state-of-the-art (SOTA) IAD methodologies. Real-IAD Variety, characterized by its unprecedented scale and diversity, redefines the benchmark standards for IAD datasets. It encompasses 160 categories spanning 28 industries, 24 material types, and 22 color variations with 198,960 images, effectively overcoming the limitations of previous datasets that were constrained by narrow category ranges and limited scenario representation.
Our comprehensive experimental analysis on Real-IAD Variety reveals several critical findings: (1) Multi-class unsupervised anomaly detection (MUAD) methods experience significant performance degradation as category count increases, with performance declining by $10–30\%$ when scaling from 30 to 160 categories; (2) Zero-shot and few-shot approaches demonstrate remarkable resilience to category scale-up, with minimal performance variation across different category counts, suggesting their superior effectiveness in handling diverse industrial scenarios; (3) Vision-language models (VLMs) leveraging external pre-trained knowledge exhibit fundamentally different scalability characteristics compared to traditional unsupervised methods, highlighting the importance of multimodal learning for large-scale IAD applications. This research underscores the pivotal role of dataset diversity and scale in advancing IAD model development. Real-IAD Variety and its associated benchmarks provide a solid foundation for future research and innovation, particularly in multi-class, multi-view and zero-/few-shot anomaly detection paradigms. We anticipate that this benchmark will facilitate the development of models with enhanced generalization capabilities, ultimately improving the reliability and efficiency of anomaly detection systems in real-world industrial applications.
Acknowledgements
This work was partially supported by the National Natural Science Foundation of China (Grant No. 62171139).